# Symmetry-Aware GFlowNets

### v1

@@ Appendix G Discussion on the Flow-Matching Objective
$\sum_{G:(G,G^{\prime})\in{\mathcal{E}}}\tilde{F}(G\rightarrow G^{\prime})=\sum_{G^{\prime\prime}:(G^{\prime},G^{\prime\prime})\in{\mathcal{E}}(G)}\tilde{F}(G^{\prime}\rightarrow G^{\prime\prime}),$
where $\tilde{F}(\cdot\rightarrow\cdot)$ denotes a permutation equivalent graph-level edge-flow function. If the rewards are given by $\tilde{R}(G)=|{\mathrm{Aut}}(G)|R(G)$ and the graph-level flow-matching condition holds for all states, then the forward policy induced by $\tilde{F}$ samples terminal states proportionally to the original reward $R$.
@@
As with the DB objective, flow-matching can be corrected at each state without scaling the reward. Outflows, $\sum_{G^{\prime\prime}:(G^{\prime},G^{\prime\prime})\in{\mathcal{E}}(G)}\tilde{F}(G^{\prime}\rightarrow G^{\prime\prime})$, are computed by summing the outputs of the neural network, which already accounts for all equivalent actions. In contrast, computing the inflows for a given graph $G^{\prime}$ involves two considerations: 1) enumerating parents of $G^{\prime}$, which may include isomorphic duplicates; 2) computing edge flows $\tilde{F}(G\rightarrow G^{\prime})$ for each $G$ such that $(G,G^{\prime})\in{\mathcal{E}}$, which do not reflect equivalence classes.
These considerations lead to the following result.
@@
where $\tilde{F}(\cdot\rightarrow\cdot)$ denotes a permutation-equivalent edge-flow function.
@@
$\sum_{s:(s,s^{\prime})\in{\mathcal{A}}}F(s\rightarrow s^{\prime})$ $=\sum_{G:(G,G^{\prime})\in{\mathcal{E}}}\frac{\text{Number of forward orbit-equivalent actions from $G$ to $G^{\prime}$}}{\text{Number of backward orbit-equivalent actions from $G^{\prime}$ to $G$}}\tilde{F}(G\rightarrow G^{\prime})$
@@
where the last equality holds in virtue of Lemma F.8. The first equality follows from the two considerations mentioned: 1) dividing by the number of backward orbit-equivalent actions to account for duplicates; 2) scaling inflows by the number of forward orbit-equivalent actions to account for duplicates. The outflows are directly computed as:
@@ Appendix H Discussion on the Fragment-based Generation
$|{\mathrm{Aut}}(G)|=1$$G$$|{\mathrm{Aut}}(G^{\prime})|=6$$G^{\prime}$$|{\mathrm{Orb}}(G,C)|=1$$|{\mathrm{Orb}}(G^{\prime},V)|=1$
@@
Number of forward orbit-equivalent actionsNumber of backward orbit-equivalent $\frac{\text{Number of forward orbit-equivalent actions}}{\text{Number of backward orbit-equivalent actions}}=\frac{|{\mathrm{Orb}}(G,V)|}{|{\mathrm{Orb}}(G\cup C,C)|}=\frac{|{\mathrm{Aut}}(G)|\cdot|{\mathrm{Aut}}(C)|}{|{\mathrm{Aut}}(G\cup C)|}.$
@@
Let $G$ represents a terminal state ($[G]\in{\mathcal{X}}$) generated by connecting k fragments $\{C_{1},\dots,C_{k}\}$. Then, the scaled rewards to offset the effects of equivalent actions are given by:
@@ Appendix I Relation to Node Orderings
Some previous work on graph generation uses a distribution over permutations (or node orderings) $\pi$, treating it as a random variable (Li et al., 2018; Chen et al., 2021). Since the node ordering determines the generation order of a graph, the joint probability over the node ordering and state is given by the following:
$P(s_{n},\pi)=P(s_{0:n},\pi)=q(\pi|s_{0:n})p_{\mathcal{A}}(s_{0:n}).$ (10)
Chen et al. (2021) derived the exact formula for $q(\pi|s_{0:n})$ and trained a model for $p_{\mathcal{A}}(s_{0:n})$. However, the joint probability $P(s_{0:n},\pi)$ can be easily obtained by multiplying graph-level transition probabilities, without needing to model $p_{\mathcal{A}}(s_{0:n})$ or adjusting for equivalent actions:
$P(s_{0:n},\pi)=p_{\mathcal{E}}(G_{0:n}).$ (11)
This result follows because, given a state sequence $s_{0:n}$, the number of different node orderings is equal to the number of possible paths generated by following equivalent actions. In other words, we can interpret different actions that are equivalent as different node orderings that induce the same state sequence. However, previous work on graph generation is not clear about how corrections for equivalent actions were made. This provides a simple formula for computing $P(s_{0:n},\pi)$ and highlights the importance of distinguishing between graphs and states. See Chen et al. (2021) for more details on using node orderings as a random variable.
@@
We trained the models for 30,000 updates using the TB objective. During the first 16,000 steps, each update used a batch of 128 trajectories, comprising 32 samples from the current policy and 96 samples drawn from the replay buffer. We used the Adam optimizer (Kingma, 2014) with the default parameters from PyTorch (Paszke et al., 2019) settings, except for the learning rates: 0.0001 for GNN layers and 0.01 for the normalizing constant $Z$. For the remaining steps, we increased batch size to 256 and annealed the learning rate to 0.00001. To encourage exploration, the policy selected actions uniformly with a probability of 0.1.

### v2

@@ Appendix G Discussion on the Flow-Matching Objective
$\sum_{G:(G,G^{\prime})\in{\mathcal{E}}}\tilde{F}(G\rightarrow G^{\prime})=\sum_{G^{\prime\prime}:(G^{\prime},G^{\prime\prime})\in{\mathcal{E}}}\tilde{F}(G^{\prime}\rightarrow G^{\prime\prime}),$
[…]
@@
As with the DB objective, FM can be corrected at each state without scaling the reward. Consider a neural network that parameterizes the edge flow function and outputs all possible outgoing edge flows $\tilde{F}(G\rightarrow G^{\prime})$ from a given graph $G$. Outflows, defined as $\sum_{G^{\prime\prime}:(G^{\prime},G^{\prime\prime})\in{\mathcal{E}}}\tilde{F}(G^{\prime}\rightarrow G^{\prime\prime})$, can be computed easily by summing the outputs of the neural network. In contrast, computing the inflows for a given graph $G^{\prime}$ involves two considerations: (1) enumerating parents of $G^{\prime}$, which may include isomorphic duplicates. To account for this, the total inflow should be divided by the number of duplicates; (2) computing edge flows $\tilde{F}(G\rightarrow G^{\prime})$ for each parent graph $G$, which may involve equivalence actions. To account for this, the total inflow should be multiplied by the number of equivalent actions. These considerations lead to the following result.
@@
[…]
@@
[…]
@@
where the last equality holds in virtue of Lemma F.8. The first equality follows from the two considerations mentioned: (1) dividing by the number of backward orbit-equivalent actions to account for duplicate parents; (2) multiplying inflows by the number of forward orbit-equivalent actions to account for duplicate actions. On the other hand, the outflows are directly computed as:
@@ Appendix H Discussion on the Fragment-based Generation
$|{\mathrm{Aut}}(G)|=1$$|{\mathrm{Aut}}(G^{\prime})|=6GG^{\prime}$$|{\mathrm{Orb}}(G,C)|=1$$|{\mathrm{Orb}}(G^{\prime},V)|=1$
@@
[…]
@@
Let $G$ represents a terminal state ($G]\in{\mathcal{X}}$) generated by connecting k fragments $\{C_{1},\dots,C_{k}\}$. Then, the scaled rewards to offset the effects of equivalent actions are given by:
@@ Appendix I Relation to Node Orderings
Some previous work on graph generation uses a distribution over permutations (or node orderings) $\pi$, treating it as a random variable (Chen et al., 2021; kong2023autoregressive; wang2025learning). If the node (or edge) ordering $\pi$ is given, the generation path $s_{0:n}=(s_{0},\dots,s_{n})$ can be determined. However, the converse is not generally true: for each state sequence $s_{0:n}$, there can be multiple compatible orderings $\pi$, complicating the computation of variational lower bound.
Chen et al. (2021) derived an exact formula for the number of node orderings consistent with a given trajectory $s_{0:n}$: it is given by the product of the sizes of the orbits encountered during the generative process, $\prod_{t=0}^{n}|{\mathrm{Orb}}(s_{t},a_{t})|$. To make this computation tractable, they approximate the orbit sizes using the color refinement algorithm. This estimate is then used to compute the joint probability $p(s_{0:n},\pi)$.
However, the joint probability can be computed easily if we consider the corresponding graph sequence $G_{0:n}$, as $p(s_{0:n},\pi)=\prod_{t=0}^{n-1}p_{\mathcal{E}}(G_{t+1}|G_{t})$. This holds because the number of different orderings that induce the same state sequence $s_{0:n}$ corresponds to the number of distinct paths generated by following transition-equivalent actions. In this view, different actions that are transition-equivalent can be interpreted as actions that induce different orderings while resulting in the same sequence of graph states. See Chen et al. (2021) for more details on using node orderings as a random variable.
@@
We trained the models for 30,000 updates using the TB objective. During the first 16,000 steps, each update used a batch of 128 trajectories, comprising 32 samples from the current policy and 96 samples drawn from the replay buffer. We used the Adam optimizer (Kingma, 2014) with the default parameters from PyTorch (Paszke et al., 2019) settings, except for the learning rates: 0.0001 for GNN layers and 0.01 for the normalizing constant $Z$. For the remaining steps, we increased batch size to 256 and annealed the learning rate to 0.00001. To encourage exploration, the policy selected actions uniformly with a probability of 0.9.